Peter is 24 years younger than his father.In 5 years' time , his father will be 3 times as old as Peter.How old is Peter now? How old will Peter's father be in 25 years' time ?
Found 2 solutions by stanbon, Vixiachase:
Answer by stanbon(75887)

Peter is 24 years younger than his father.
p = f - 24
--------------------------
In 5 years' time , his father will be 3 times as old as Peter
f+5 = 3(p+5)
------------------------
How old is Peter now?
Substitute for "f" and solve for "p"::
p+24 + 5 = 3p + 15
p+29 = 3p+15
2p = 14
Peter's age = 7 yrs
--------------------------
How old will Peter's father be in 25 years' time ?
Now:: f = p+24 = 31 yrs
Ans: 31+25 = 56 yrs
--------------
Cheers,
Stan H.
-----------

Answer by Vixiachase(22)